OpenStudy (anonymous):

solve for n: Square root--n+5--+n-1 not sure how to type the square root sign but n+5 is under it..

OpenStudy (anonymous):

is it \[\sqrt{n-5} = n+1\]

OpenStudy (anonymous):

or \[\sqrt{n-5} = n-1\]

OpenStudy (anonymous):

its n+5= n-1 with square root n+5

OpenStudy (anonymous):

u can pose this question as : squareroot(n-5) = n-1 and then people will understand ans is n = 4 and n= -1 do u need the working ??